# Six-Point Method for Multi-Camera Systems with Reduced Solution Space

### v1

@@ Abstract
Relative pose estimation using point correspondences (PC) is a widely used technique. A minimal configuration of six PCs is required for generalized cameras. In this paper, we present several minimal solvers that use six PCs to compute the 6DOF relative pose of a multi-camera system, including a minimal solver for the generalized camera and two minimal solvers for the practical configuration of two-camera rigs. The equation construction is based on the decoupling of rotation and translation. Rotation is represented by Cayley or quaternion parametrization, and translation can be eliminated by using the hidden variable technique. Ray bundle constraints are found and proven when a subset of PCs relate the same cameras across two views. This is the key to reducing the number of solutions and generating numerically stable solvers. Moreover, all configurations of six-point problems for multi-camera systems are enumerated. Extensive experiments demonstrate that our solvers are more accurate than the state-of-the-art six-point methods, while achieving better performance in efficiency.
@@ 1 Introduction
Camera models have a significant impact on computing the relative pose. A pinhole or perspective camera model is used to model a single camera [19], and more complicated cameras like multi-camera systems require a generalized camera model for modeling [12, 42, 48]. A generalized camera can be formed by abstracting the landmark observations into spatial rays, which do not necessarily require originating from the projection center. This paper is principally concerned with a multi-camera setup comprising several cameras that have been installed rigidly. As we shall demonstrate in this paper, $n$ point correspondences (PCs) for a generalized camera can be represented similarly by using $2n$ single cameras in a multi-camera system. It is a common belief that the standard epipolar geometry using five PCs is incapable of recovering the scale of translation [19]. Conversely, the translation scale of multi-camera systems can be uniquely determined, and the minimum requirement for solving the relative pose increases from five to six PCs across both views [46].
Due to the presence of outliers in PCs, a robust estimator is essential for accurate relative pose estimation and outlier rejection. The random sample consensus (RANSAC) framework [9] and its various adaptations [30, 44, 1, 3] are widely employed in computer vision community. A minimal solver is the core component in the RANSAC framework. Based on the epipolar geometry of each PC, one constraint for solving relative pose can be established [19]. Various methods for computing relative pose of a single camera are known as five-point methods [40, 47, 34, 26, 27, 8]. The six-point method [46] is the first minimal solver proposed for computing relative pose of a multi-camera system. Numerous methods have been introduced subsequently, such as the enhanced version of the six-point method [5], the seventeen-point linear solvers [35, 22], an iterative optimization-based solver [24], and a global optimization-based solver [53].
This paper estimates the full DOF relative pose for a multi-camera system using PCs. The main contributions of this work are as follows:
  • We estimate 6DOF relative pose from a minimal number of six PCs for multi-camera systems. A generic minimal solver for the generalized camera and two minimal solvers for popular configurations of two-camera rigs are proposed. Moreover, we enumerate all configurations of minimal six-point problems for multi-camera systems.
@@
  • We exploit a framework by decoupling rotation and translation. Rotation can be represented by Cayley or quaternion parametrization, and translation is eliminated by using the hidden variable technique. This framework can easily deal with relative pose estimation for other configurations of multi-camera systems.
@@ 2 Related Work
The research on relative pose estimation remains active in geometric vision, with a lot of classical solvers existing in this area. First, these solvers can be divided into the relative pose estimation for single cameras [19, 40, 47, 34, 26, 27, 8], multi-camera systems or generalized cameras [46, 35, 22, 24, 53]. These cameras can be calibrated or partially uncalibrated with unknown focal length or radial distortion.
Second, the relative pose estimation solvers can be categorized as minimal solvers [40, 46], non-minimal solvers [22, 25, 24, 53, 51] and linear solvers [20, 35]. The minimal solvers aim to estimate relative pose using the minimum number of geometric primitives. The non-minimal solvers leverage all the feature correspondences to compute the relative pose. The linear solvers require more feature correspondences than the minimal solvers, and their primary purpose is to provide a straightforward and efficient solution. To achieve computational efficiency, the linear solvers often ignore implicit constraints on unknown parameters. Conversely, the minimal and non-minimal solvers usually exploit implicit constraints while solving for unknowns. When dealing with feature correspondences that may contain outliers, minimal solvers play a crucial role in ensuring robust estimation. For instance, RANSAC and its variants depend heavily on efficient minimal solvers.
The combination of the two aspects mentioned above produces many subcategories. In this paper, we focus on exploring the minimal solvers for multi-camera systems using PCs. A minimal solver was first proposed to solve the relative pose of multi-camera systems using $6$ PCs [46]. Then, a linear solver taking $17$ PCs was introduced in [35, 22]. Several solvers were developed for applying to structure-from-motion with special configurations [54, 21]. Several non-minimal solvers were proposed which leverage either local optimization [24] or global optimization [53] to determine optimal relative poses. Some solvers required the motion priors of multi-camera systems, including known rotation axis [32, 49, 37] and Ackermann motion [31]. Moreover, an efficient solver was achieved by implementing a first-order approximation of relative pose  [50].
@@ 3.1 Geometric Constraints
Figure 1: Relative pose estimation for a multi-camera system. A point is captured by perspective camera $C_{i}$ in view 1 and camera $C_{i^{\prime}}$ in view 2. $\{\mathbf{Q}_{i},\mathbf{s}_{i}\}$ and $\{\mathbf{Q}_{i^{\prime}},\mathbf{s}_{i^{\prime}}\}$ are extrinsic parameters for $C_{i}$ and $C_{i^{\prime}}$, respectively. The related point correspondence can be described by the two-view geometry of cameras $C_{i}$ and $C_{i^{\prime}}$.
Fig. 1 illustrates a multi-camera system consisting of multiple perspective cameras, assuming that the intrinsic and extrinsic parameters of perspective cameras have been calibrated. The extrinsic parameters of the $i$-th camera are denoted as $\{\mathbf{Q}_{i},\mathbf{s}_{i}\}$, which are the relative rotation and translation to the reference of the multi-camera system. Let $\mathbf{R}$ and $\mathbf{t}$ denote the relative rotation and translation between the first and second views, respectively.
A PC in a multi-camera system establishes the relationship of a point captured by two perspective cameras across two different views. Let the $k$-th PC be denoted by $(\mathbf{x}_{k},\mathbf{x}^{\prime}_{k},i_{k},i^{\prime}_{k})$. This indicates that the $i_{k}$-th camera captures a point in the first view, which is represented by its homogeneous coordinate in the normalized image plane denoted as $\mathbf{x}_{k}$. Furthermore, this same point is also captured by the $i^{\prime}_{k}$-th camera in the second view, which is represented by its homogeneous coordinate as $\mathbf{x}^{\prime}_{k}$. In the following text, we omit the subscript $k$ of camera indices $i$ and $i^{\prime}$ to simplify the notation. In a multi-camera system, the essential matrices for different PCs are typically distinct, which sets it apart from the case of two-view geometry for a single camera. Thus, one constraint of epipolar geometry [19] introduced by the $k$-th PC is
$\mathbf{x}_{k}^{\prime T}\mathbf{E}_{k}\mathbf{x}_{k}=0,$ (1)
@@
Here $\{\mathbf{R}_{ii^{\prime}},\mathbf{t}_{ii^{\prime}}\}$ is the relative rotation and translation from camera $i$ in the first view to camera $i^{\prime}$ in the second view. According to Fig. 1, it is obtained by a composition of three spatial transformations as
$\begin{bmatrix}\mathbf{R}_{ii^{\prime}}&{\mathbf{t}_{ii^{\prime}}}\\ {{\mathbf{0}}}&{1}\\ \end{bmatrix}=$ $\begin{bmatrix}{\mathbf{Q}_{i^{\prime}}}&{\mathbf{s}_{i^{\prime}}}\\ {{\mathbf{0}}}&{1}\\ \end{bmatrix}^{-1}\begin{bmatrix}\mathbf{R}&{\mathbf{t}}\\ {{\mathbf{0}}}&{1}\\ \end{bmatrix}\begin{bmatrix}{\mathbf{Q}_{i}}&{\mathbf{s}_{i}}\\ {{\mathbf{0}}}&{1}\\ \end{bmatrix}$ (11)
@@
Based on the above equation, it can be seen that Eq. (1) is bilinear in the relative pose $(\mathbf{R},\mathbf{t})$.
@@ 3.2 Relative Pose Parameterization
We need to parametrize the relative pose of multi-camera systems. Rotation can be parameterized by Cayley parametrization, quaternions, Euler angles, direction cosine matrix (DCM), etc. Cayley and quaternion parameterizations have shown superiority in minimal problems [52]. Rotation $\mathbf{R}$ using Cayley parameterization can be written as
$\mathbf{R}_{\text{cayl}}=\small{\frac{1}{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1}}\ .$
$\small{\begin{bmatrix}{1+q_{x}^{2}-q_{y}^{2}-q_{z}^{2}}&2q_{x}q_{y}-2q_{z}&2q_{x}q_{z}+2q_{y}\\ 2q_{x}q_{y}+2q_{z}&1-q_{x}^{2}+q_{y}^{2}-q_{z}^{2}&2q_{y}q_{z}-2q_{x}\\ 2q_{x}q_{z}-2q_{y}&2q_{y}q_{z}+2q_{x}&{1-q_{x}^{2}-q_{y}^{2}+q_{z}^{2}}\end{bmatrix}},$ (19)
@@
$\mathbf{R}_{\text{quat}}=$
$\small{\begin{bmatrix}q_{w}^{2}+q_{x}^{2}-q_{y}^{2}-q_{z}^{2}&2q_{x}q_{y}-2q_{w}q_{z}&2q_{x}q_{z}+2q_{w}q_{y}\\ 2q_{x}q_{y}+2q_{w}q_{z}&q_{w}^{2}-q_{x}^{2}+q_{y}^{2}-q_{z}^{2}&2q_{y}q_{z}-2q_{w}q_{x}\\ 2q_{x}q_{z}-2q_{w}q_{y}&2q_{y}q_{z}+2q_{w}q_{x}&q_{w}^{2}-q_{x}^{2}-q_{y}^{2}+q_{z}^{2}\end{bmatrix}},$ (23)
where $[q_{w},q_{x},q_{y},q_{z}]^{T}$ is a normalized quaternion vector satisfying the following constraints:
$q_{w}^{2}+q_{x}^{2}+q_{y}^{2}+q_{z}^{2}=1.$ (24)
Note that $180^{\circ}$ rotations are prohibited in Cayley parameterization, but this is a rare case for usual image pairs. In practice, it has been widely used in minimal problems [45, 24, 54, 52]. By contrast, quaternion parameterization does not have any degeneracy. However, quaternion introduces more variables than Cayley, so its solvers are usually less efficient. We construct solvers using both of these two parameterizations for completeness. In the following, we introduce the solver generation procedure based on Cayley parameterization. This procedure can also be applied to quaternion parameterization straightforwardly.
@@ 3.3 Equation System Construction
where the entries of ${\mathbf{M}}$ are quadratic polynomials with three unknowns $q_{x},q_{y},q_{z}$. The $i$-th row corresponds to the constraint of $i$-th PC. It can be seen that ${\mathbf{M}}$ has a null vector. Thus, the determinants of all the $4\times 4$ submatrices of ${\mathbf{M}}$ should be zero.
Moreover, when a subset of PCs relates the same perspective cameras across two views, there is a property for this scenario. This scenario can be used to reduce the number of solutions and generate numerically stable solvers.
@@
Denote $\mathcal{S}$ as a matrix set with elements satisfying $\mathbf{N}={\mathbf{M}}([k_{1},k_{2},k_{3}],1:3)$, where $\mathbf{N}$ is formed from rows $\{k_{1},k_{2},k_{3}\}$ and columns $\{1,2,3\}$ of $\mathbf{M}$. In addition, $k_{1}$-th, $k_{2}$-th, and $k_{3}$-th PCs are captured by the same perspective cameras across two views and $k_{1}<k_{2}<k_{3}$. Then $\operatorname*{rank}(\mathbf{N})=2$, $\forall\mathbf{N}\in\mathcal{S}$ holds for non-degenerate cases.
@@
Let us investigate an arbitrary element of $\mathcal{S}$. Denote it is the $k$-th element $\mathbf{N}_{k}$ in $\mathcal{S}$. The extrinsic parameters of the related perspective camera in view 1 is $\{\mathbf{Q}_{i},\mathbf{s}_{i}\}$, and the extrinsic parameters of the related perspective camera in view 2 is $\{\mathbf{Q}_{i^{\prime}},\mathbf{s}_{i^{\prime}}\}$. First we prove that $\operatorname*{rank}(\mathbf{N}_{k})\leq 2$. To achieve this goal, we need to prove that the null space of $\mathbf{N}_{k}$ is not empty. Since $k_{1}$-th, $k_{2}$-th, and $k_{3}$-th PCs are captured by the same perspective cameras across two views, their related essential matrices are the same. According to Eq. (15), the essential matrix is
@@
$\frac{1}{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1}\mathbf{A}_{k}\bar{\mathbf{t}}=\mathbf{0},$ (34)
$\Rightarrow$ $\mathbf{A}_{k}(\mathbf{t}+\mathbf{R}\mathbf{s}_{i}-\mathbf{s}_{i^{\prime}})=\mathbf{0},$ (35)
@@
Substituting this equation into Eq. (34), we can see that the null space of $\mathbf{N}_{k}$ is not empty. Next we prove that $\operatorname*{rank}(\mathbf{N}_{k})\geq 2$. We achieve this goal using proof by contradiction. If $\operatorname*{rank}(\mathbf{N}_{k})\leq 1$, then $\operatorname*{rank}({\mathbf{M}}([k_{1},k_{2},k_{3}],{1:4}))\leq 2$ considering that ${\mathbf{M}}([k_{1},k_{2},k_{3}],1:4)$ has one more column than $\mathbf{N}_{k}$. This means the three PCs provide at most two independent constraints for the relative pose estimation. This does not hold true for non-degenerate cases, hence invalidating the assumption that $\operatorname*{rank}(\mathbf{N}_{k})\leq 1$. It can be seen that the rank of $\mathbf{N}$ is 2 in non-degenerate cases. ∎
We call the constraints in Theorem 1 as the ray bundle constraints. In computer graphics, a ray bundle is a collection of light rays that share a common origin and propagate in different directions. In addition, a factor $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$ can be factored out to simplify the equation system. Please see the supplementary material for proof. It will generate more efficient solvers and sometimes avoid extraneous roots. In summary, the whole polynomial equation system consists of two types of constraints, which can be written as
${\mathcal{E}}_{1}\triangleq\{\operatorname*{quot}(\det(\mathbf{N}),q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1)=0\ |$
$\mathbf{N}\in 4\times 4\text{ submatrices of }{\mathbf{M}}\},$ (41)
@@
where $\operatorname*{quot}(a,b)$ means quotient of $a$ divided by $b$, and $\det(\cdot)$ means the determinant operator.
There are $15$ equations of degree $6$ and some equations of degree 4 in ${\mathcal{E}}_{1}$ and ${\mathcal{E}}_{2}$, respectively. Note that the number of equations in ${\mathcal{E}}_{2}$ varies for different PC configurations. ${\mathcal{E}}_{2}$ can be an empty set for certain PC configurations of multi-camera systems.
Once the rotation parameters $\{q_{x},q_{y},q_{z}\}$ have been obtained, the translation $[t_{x},t_{y},t_{z}]^{T}$ can be recovered by first calculating a vector in the null space of $\mathbf{M}$, and then normalizing the vector by dividing its last entry.

### v2

# {textblock}

@@ Abstract
Relative pose estimation using point correspondences (PC) is a widely used technique. A minimal configuration of six PCs is required for two views of generalized cameras. In this paper, we present several minimal solvers that use six PCs to compute the 6DOF relative pose of multi-camera systems, including a minimal solver for the generalized camera and two minimal solvers for the practical configuration of two-camera rigs. The equation construction is based on the decoupling of rotation and translation. Rotation is represented by Cayley or quaternion parametrization, and translation can be eliminated by using the hidden variable technique. Ray bundle constraints are found and proven when a subset of PCs relate the same cameras across two views. This is the key to reducing the number of solutions and generating numerically stable solvers. Moreover, all configurations of six-point problems for multi-camera systems are enumerated. Extensive experiments demonstrate the superior accuracy and efficiency of our solvers compared to state-of-the-art six-point methods. The code is available at https://github.com/jizhaox/relpose-6pt.
Keywords:
Minimal solver Relative pose estimation Point correspondence Multi-camera system Ray bundle constraint
@@ 1 Introduction
Camera models have a significant impact on computing the relative pose. A pinhole or perspective camera model is used to model a single camera [19], and more complicated cameras like multi-camera systems necessitate the use of a generalized camera model [12, 49]. A generalized camera encapsulates various imaging systems by representing the landmark observations as spatial rays, which do not necessarily require originating from the projection center [43]. This paper is principally concerned with a multi-camera system comprising several cameras that have been installed rigidly. As we shall demonstrate in this paper, $n$ point correspondences (PCs) for a generalized camera can be represented similarly by using $2n$ single cameras in a multi-camera system. It is a proven fact that the standard epipolar geometry using five PCs is incapable of recovering the scale of translation [19]. Conversely, the translation scale of multi-camera systems can be uniquely determined, and the minimum requirement for solving the relative pose increases from five to six PCs across both views [48].
Due to the presence of outliers in PCs, a robust estimator is essential for accurate relative pose estimation and outlier rejection. The random sample consensus (RANSAC) framework [9] and its various adaptations [31, 45, 1, 2] are widely employed in computer vision community. A minimal solver is the core component in the RANSAC framework. Using the epipolar geometry corresponding to each PC, a constraint can be derived to solve for the relative pose [19]. Various methods for computing the relative pose of a single camera are known as five-point methods [42, 46, 35, 26, 27, 8]. The six-point method [48] is the first minimal solver proposed for computing the relative pose of a multi-camera system. Numerous methods have been introduced subsequently, such as the enhanced version of the six-point method [5], the seventeen-point linear solvers [36, 22], an iterative optimization-based solver [24], and a global optimization-based solver [55].
This paper utilizes PCs to estimate the full DOF relative pose for multi-camera systems. The main contributions of this work are summarized as follows:
  • We estimate 6DOF relative pose from a minimal number of six PCs for multi-camera systems. By decoupling rotation and translation, a generic minimal solver for the generalized camera and two minimal solvers for popular configurations of two-camera rigs are proposed.
@@
  • All configurations of minimal six-point problems for multi-camera systems are first enumerated using graph enumeration and the Pólya enumeration theorem. Totally there are $5953$ cases. Moreover, we enumerate all the distinct graphs in a recursive way with 50 lines of Matlab code only.
@@ 2 Related Work
The research on relative pose estimation remains active in geometric vision, with a lot of classical solvers existing in this area. First, these solvers can be divided into the relative pose estimation for single cameras [19, 42, 46, 35, 26, 27, 8] and generalized cameras [48, 36, 22, 24, 55]. These cameras can be calibrated or partially uncalibrated with unknown focal length or radial distortion.
Second, the relative pose estimation solvers can be categorized as minimal solvers [42, 48], non-minimal solvers [22, 25, 24, 55, 52] and linear solvers [20, 36]. The minimal solvers aim to estimate relative pose using the minimum number of geometric primitives. The non-minimal solvers leverage all the feature correspondences to compute the relative pose. The linear solvers demand a higher number of feature correspondences compared to minimal solvers and provide a straightforward and efficient solution. To achieve computational efficiency, the linear solvers often ignore implicit constraints on unknown parameters. Conversely, the minimal and non-minimal solvers usually exploit implicit constraints while solving for unknown parameters. When dealing with feature correspondences that may contain outliers, minimal solvers play a crucial role in ensuring robust estimation. For instance, RANSAC and its variants depend heavily on efficient minimal solvers.
The integration of these two aspects yields numerous subcategories. This paper specifically investigates minimal solvers for multi-camera systems using PCs. A minimal solver with $64$ solutions was first proposed to address the relative pose of multi-camera systems using $6$ PCs [48]. They also propose a special case with $56$ solutions when the extrinsic parameters satisfy a certain condition [48]. Then, a linear solver taking $17$ PCs was introduced in [36, 22]. Several solvers were proposed for applying to structure-from-motion with special configurations [56, 21]. Moreover, a few cases for semi-generalized cameras were proposed in [56], which is a rather small subset of our enumerated cases. Several non-minimal solvers were proposed which leverage either local optimization [24] or global optimization [55] to determine optimal relative poses. Some solvers required the motion priors of multi-camera systems, including known rotation axis [33, 50, 38] and Ackermann motion [32]. Moreover, an efficient solver was achieved by implementing a first-order approximation of relative rotation [51]. By exploiting the additional affine parameters besides PCs, a minimal solver with two affine correspondences was proposed in [14]. Three significant differences between [14] and our method are clarified in the supplementary material.
@@ 3.1 Geometric Constraints
Fig. 1 illustrates a multi-camera system consisting of multiple perspective cameras, assuming that both the intrinsic and extrinsic parameters of these cameras have been calibrated. The extrinsic parameters of the camera $C_{i}$ are represented as $\{\mathbf{Q}_{i},\mathbf{s}_{i}\}$, where $\mathbf{Q}_{i}$ represents the rotation and $\mathbf{s}_{i}$ represents the translation with respect to the multi-camera system’s reference. Let $\mathbf{R}$ denote the rotation and $\mathbf{t}$ denote the translation between the first and second views.
Figure 1: Relative pose estimation for a multi-camera system. A point is observed by perspective camera $C_{i}$ in view 1 and by camera $C_{i^{\prime}}$ in view 2. $\{\mathbf{Q}_{i},\mathbf{s}_{i}\}$ and $\{\mathbf{Q}_{i^{\prime}},\mathbf{s}_{i^{\prime}}\}$ are extrinsic parameters for $C_{i}$ and $C_{i^{\prime}}$, respectively. The related point correspondence is described by two-view epipolar geometry of cameras $C_{i}$ and $C_{i^{\prime}}$.
A PC in a multi-camera system establishes the relationship of a point captured by two cameras across different views. Let the $k$-th PC be denoted by $(\mathbf{x}_{k},\mathbf{x}^{\prime}_{k},i_{k},i^{\prime}_{k})$. This indicates that the $i_{k}$-th camera observes a point in view 1, which is represented by its homogeneous coordinate $\mathbf{x}_{k}$ in the normalized image plane. Furthermore, this same point is also observed by the $i^{\prime}_{k}$-th camera in view 2, which is represented by its homogeneous coordinate as $\mathbf{x}^{\prime}_{k}$. For simplicity, we omit the subscript $k$ from camera indices $i$ and $i^{\prime}$ to simplify the notation. In a multi-camera system, essential matrices vary for different PCs, distinguishing them from the two-view epipolar geometry of single cameras. Therefore, one constraint of epipolar geometry [19] induced by the $k$-th PC is
$\mathbf{x}_{k^{\prime}}^{T}\mathbf{E}_{k}\mathbf{x}_{k}=0,$ (1)
@@
Here $\{\mathbf{R}_{ii^{\prime}},\mathbf{t}_{ii^{\prime}}\}$ are the relative rotation and translation from camera $i$ in the first view to camera $i^{\prime}$ in the second view. According to Fig. 1, it is obtained by a composition of three spatial transformations as
@@
Based on the above equation, it can be seen that Eq. (1) is bilinear in the relative pose $\{\mathbf{R},\mathbf{t}\}$.
@@ 3.2 Relative Pose Parameterization
We need to parametrize the relative pose of multi-camera systems. Rotation can be parameterized using Cayley parametrization, quaternions, Euler angles, direction cosine matrix (DCM), etc. Cayley and quaternion parameterizations have demonstrated superiority in minimal problems [54]. Rotation matrix $\mathbf{R}$ using Cayley parameterization can be expressed as
$\mathbf{R}_{\text{cayl}}=\small{\frac{1}{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1}}{\begin{bmatrix}{1+q_{x}^{2}-q_{y}^{2}-q_{z}^{2}}&2q_{x}q_{y}-2q_{z}&2q_{x}q_{z}+2q_{y}\\ 2q_{x}q_{y}+2q_{z}&1-q_{x}^{2}+q_{y}^{2}-q_{z}^{2}&2q_{y}q_{z}-2q_{x}\\ 2q_{x}q_{z}-2q_{y}&2q_{y}q_{z}+2q_{x}&{1-q_{x}^{2}-q_{y}^{2}+q_{z}^{2}}\end{bmatrix}},$ (5)
@@
$\mathbf{R}_{\text{quat}}={\begin{bmatrix}q_{w}^{2}+q_{x}^{2}-q_{y}^{2}-q_{z}^{2}&2q_{x}q_{y}-2q_{w}q_{z}&2q_{x}q_{z}+2q_{w}q_{y}\\ 2q_{x}q_{y}+2q_{w}q_{z}&q_{w}^{2}-q_{x}^{2}+q_{y}^{2}-q_{z}^{2}&2q_{y}q_{z}-2q_{w}q_{x}\\ 2q_{x}q_{z}-2q_{w}q_{y}&2q_{y}q_{z}+2q_{w}q_{x}&q_{w}^{2}-q_{x}^{2}-q_{y}^{2}+q_{z}^{2}\end{bmatrix}},$ (6)
where $[q_{w},q_{x},q_{y},q_{z}]^{T}$ is a quaternion vector satisfying the normalization constraint $q_{w}^{2}+q_{x}^{2}+q_{y}^{2}+q_{z}^{2}=1$.
Note that $180^{\circ}$ rotations are not allowed in Cayley parameterization, although this is uncommon for typical image pairs. In practical applications, Cayley parameterization has been extensively utilized in minimal problems [47, 24, 56, 54]. By contrast, quaternion parameterization does not have any degeneracy. However, quaternion introduces more variables than Cayley, so its solvers are usually less efficient. We construct solvers using both of these two parameterizations for completeness. In the following, we introduce the solver generation procedure based on Cayley parameterization. This approach can also be directly extended to quaternion parameterization.
@@ 3.3 Equation System Construction
where the entries of ${\mathbf{M}}$ are quadratic polynomials in three unknowns $q_{x},q_{y},q_{z}$. The $i$-th row corresponds to the constraint associated with the $i$-th PC. It can be observed that ${\mathbf{M}}$ has a null vector. Hence, the determinants of all the $4\times 4$ submatrices of ${\mathbf{M}}$ must be zero.
Moreover, when a subset of PCs relates to the same perspective cameras across two views, there exists a specific property for this scenario. This property can be used to reduce the number of solutions and generate numerically stable solvers. Taking the number of a subset of PCs equal to 3 as an introductory example. When the number of a subset of PCs is greater than or equal to 3, it can be directly derived from Theorem 1.
@@
Denote $\mathcal{S}$ as a matrix set with elements satisfying $\mathbf{N}={\mathbf{M}}([k_{1},k_{2},\\ k_{3}],1:3)$, where $\mathbf{N}$ is formed from rows $\{k_{1},k_{2},k_{3}\}$ and columns $\{1,2,3\}$ of $\mathbf{M}$. In addition, $k_{1}$-th, $k_{2}$-th, and $k_{3}$-th PCs are captured by the same perspective camera in each view and $k_{1}<k_{2}<k_{3}$. Then $\operatorname*{rank}(\mathbf{N})=2$, $\forall\mathbf{N}\in\mathcal{S}$ holds for non-degenerate cases.
@@
Let’s proceed with the investigation of an arbitrary element in $\mathcal{S}$. We denote $\mathbf{N}_{k}$ as the $k$-th element in $\mathcal{S}$. The extrinsic parameters of the corresponding perspective camera in view 1 are denoted by $\{\mathbf{Q}_{i},\mathbf{s}_{i}\}$, and for view 2, they are denoted by $\{\mathbf{Q}_{i^{\prime}},\mathbf{s}_{i^{\prime}}\}$.
We begin by proving that $\operatorname*{rank}(\mathbf{N}_{k})\leq 2$. To accomplish this objective, we must prove that the null space of $\mathbf{N}_{k}$ is non-empty. Given that the $k_{1}$-th, $k_{2}$-th, and $k_{3}$-th PCs are observed by the same perspective camera in each view, their associated essential matrices remain unchanged. Referring to Eq. (4), the essential matrix can be parametrized as
@@
$\frac{1}{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1}\mathbf{A}_{k}\bar{\mathbf{t}}=\mathbf{0}$ $\Rightarrow\mathbf{A}_{k}(\mathbf{t}+\mathbf{R}\mathbf{s}_{i}-\mathbf{s}_{i^{\prime}})=\mathbf{0},$ (11)
@@
Substituting this equation into Eq. (11), we can observe that the null space of $\mathbf{N}_{k}$ is indeed not empty.
Next we aim to prove that $\operatorname*{rank}(\mathbf{N}_{k})\geq 2$. This goal can be achieved using proof by contradiction. If $\operatorname*{rank}(\mathbf{N}_{k})\leq 1$, then $\operatorname*{rank}({\mathbf{M}}([k_{1},k_{2},k_{3}],{1:4}))\leq 2$ given that ${\mathbf{M}}([k_{1},k_{2},k_{3}],1:4)$ includes an extra column compared to $\mathbf{N}_{k}$. This degenerate case implies that the three PCs offer no more than two independent constraints for the relative pose estimation. This does not hold true for non-degenerate cases, hence invalidating the assumption that $\operatorname*{rank}(\mathbf{N}_{k})\leq 1$. It can be seen that the rank of $\mathbf{N}$ is 2 in non-degenerate cases. ∎
We call the constraints in Theorem 1 as the ray bundle constraints. In computer graphics, a ray bundle is a collection of light rays that share a common origin and propagate in different directions. In addition, a factor $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$ can be factored out to simplify the equation system. Please see the supplementary material for proof. It leads to the generation of more efficient solvers while also potentially avoiding extraneous roots. In summary, the whole polynomial equation system consists of two types of constraints, which can be written as
${\mathcal{E}}_{1}\triangleq\{\operatorname*{quot}(\det(\mathbf{N}),q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1)=0\ |\ \mathbf{N}\in\small{4\times 4\text{ submatrices of }}{\mathbf{M}}\},$ (14)
@@
where $\operatorname*{quot}(a,b)$ denotes the division quotient of $a$ by $b$, and $\det(\cdot)$ represents the determinant operator.
In ${\mathcal{E}}_{1}$, there are $15$ equations of degree $6$, while ${\mathcal{E}}_{2}$ contains some equations of degree 4. It should be noted that the number of equations in ${\mathcal{E}}_{2}$ varies depending on the specific configurations of the PCs. For certain PC configurations of multi-camera systems, ${\mathcal{E}}_{2}$ may be empty.
After obtaining the rotation parameters $\{q_{x},q_{y},q_{z}\}$, the translation $[t_{x},t_{y},t_{z}]^{T}$ can be determined by initially calculating a vector within the null space of $\mathbf{M}$, followed by normalization where the last entry of the vector is divided.